# LaCo: Large Language Model Pruning via Layer Collapse

### v1

@@ Abstract
Large language models (LLMs) based on transformer are witnessing a notable trend of size expansion, which brings considerable costs to both model training and inference. However, existing methods such as model quantization, knowledge distillation, and model pruning are constrained by various issues, including hardware support limitations, the need for extensive training, and alterations to the internal structure of the model. In this paper, we propose a concise layer-wise pruning method called Layer Collapse (LaCo), in which rear model layers collapse into a prior layer, enabling a rapid reduction in model size while preserving the model structure. Comprehensive experiments show that our method maintains an average task performance of over 80% at pruning ratios of 25-30%, significantly outperforming existing state-of-the-art structured pruning methods. We also conduct post-training experiments to confirm that the proposed pruning method effectively inherits the parameters of the original model. Finally, we discuss our motivation from the perspective of layer-wise similarity and evaluate the performance of the pruned LLMs across various pruning ratios.
LaCo: Large Language Model Pruning via Layer Collapse
[…]
@@ 1 Introduction
Recently, large language models (LLMs) based on transformer Vaswani et al. (2017) have showcased impressive capabilities across diverse tasks. However, the prevailing trend in model development leans towards larger scales, placing substantial demands on computational power and resulting in challenges for many researchers and companies.
To mitigate the above challenge, various approaches have been explored to reduce the inference and training costs of models or to derive a smaller model from an LLM, including model quantization Dettmers et al. (2022); Yao et al. (2022); Xiao et al. (2023), knowledge distillation Liu et al. (2022); Hsieh et al. (2023); Shridhar et al. (2023), and model pruning Zhang et al. (2022); Frantar and Alistarh (2023); Ma et al. (2023). However, existing solutions exhibit certain notable drawbacks. Model quantization significantly impacts model performance and typically necessitates specific hardware support. Knowledge distillation often requires retraining a smaller model at a higher training cost and is usually task-specific. Existing Model pruning can be categorized into non-structured and structured pruning. Non-structured pruning often involves model sparsity, which generally leads to notable performance loss and also relies on hardware support. Structured pruning entails removing specific modules from the model, but it frequently modifies the model structure, diminishing the portability of the model.
@@
Specifically, we observe that for certain layers, merging the differential values of the parameters of this layer and the consecutive layers following it often does not result in significant model performance loss, as illustrated in Figure 1.
We term it the Reserving-Differences-while-Seeking-Common (RDSC) Layer Merge, as it incorporates parameter differencing and merging. Building upon this insight, we introduce a streamlined yet potent layer-wise pruning method dubbed Layer Collapse (LaCo), in which rear layers collapse into a prior layer, with the objective of preserving the model’s output representation as closely as feasible to the representation before pruning. In this paper:
@@
$\bullet$ The Layer Collapse also possesses excellent properties of preserving the internal structure of model layers, such as maintaining intermediate dimensions unchanged. Hence, it can be quickly adapted to existing applications.
$\bullet$ We further conduct post-training on the pruned model, confirming that the Layer Collapse can efficiently inherit parameters, requiring only a minimal amount of training to restore the pruned model to the loss convergence level of the original model. Furthermore, we discuss our motivation and evaluate the performance of pruned models using LaCo across different pruning ratios.
@@ 2.1 Reserving-Differences-while-Seeking-Common Layer Merge
Given an LLM, for its $l$-th layer, we denote all its parameters, including those in self-attention (SAN) and MLP as $\boldsymbol{\theta}_{l}$. For the $m$ consecutive layers following it, we merge the parameters of $\boldsymbol{\theta_{l+1}},\boldsymbol{\theta_{l+2}},\cdots,\boldsymbol{\theta_{l+m}}$ into $\boldsymbol{\theta_{l}}$ to form $\boldsymbol{\theta^{*}_{l}}$:
$\boldsymbol{\theta}^{*}_{l}$ $=\boldsymbol{\theta}_{l}+(\boldsymbol{\theta}_{l+1}-\boldsymbol{\theta}_{l})+\cdots+(\boldsymbol{\theta}_{i+m}-\boldsymbol{\theta}_{l})$ (1)
$=\boldsymbol{\theta}_{l}+\sum_{k=1}^{m}(\boldsymbol{\theta}_{l+k}-\boldsymbol{\theta}_{l})$
where $(\boldsymbol{\theta_{l+k}}-\boldsymbol{\theta_{l})}$ is the parameter difference between each subsequent layer and the $l$-th layer. During parameter differencing and merging, given identical layer structures, we independently apply these processes to both SAN and MLP. After that, these $m$ consecutive layers will be discarded. Subsequent model pruning will continuously involve RDSC Layer Merge, a process that can be regarded as the continual collapse of layers onto specific layers, hence the name Layer Collapse.
@@
Algorithm 1 Layer Collapse
1:
2:LLM $\mathcal{M}$
3:Number of layers combined in each merge $\mathcal{C}$
4:Layer range considered $[\mathcal{L},\mathcal{H}]$
5:Minimum interval between two adjacent merged layers $\mathcal{I}$
6:Few-shot Calibration Samples $\mathcal{D}$
7:Threshold for representation similarity $\mathcal{T}$
8:Pruned LLM $\mathcal{M}^{*}$
@@
10:$l\leftarrow\mathcal{H}-\mathcal{C}$
11:while $l>=\mathcal{L}$ do
12:     $\mathcal{K}\leftarrow$ Min($\mathcal{C}-1$, $\mathcal{M}^{*}$)$-l$)
13:     $\mathcal{M}_{tmp}\leftarrow$ $\mathcal{M}^{*}$, $l$, $\mathcal{K}$)
14:     $s\leftarrow$ Avg_Cos_Sim($\mathcal{M}_{tmp}$, $\mathcal{M}$, $\mathcal{D}$)
15:     if $s>\mathcal{T}$ then
16:         *$\mathcal{M}^{*}\leftarrow\mathcal{M}_{tmp}$
17:         $l\leftarrow l-\mathcal{I}$
18:         if $l>$ $\mathcal{M}^{*}$then
19:              $l\leftarrow$$\mathcal{M}^{*}$)$-\mathcal{C}$
20:         end if
21:     else
22:         $l\leftarrow l-1$
23:     end if
24:end while
25:return $\mathcal{M}^{*}$
@@ 2.2 Layer Collapse
We dynamically merge adjacent layers starting from the topmost layer of the model, while ensuring that the output representation of the pruned model on few-shot calibration samples remains as similar as possible to the original model, to ensure minimal performance loss. Algorithm 1 summarizes the procedure of Layer Collapse:
(1) Preparation (line 1-6)
@@
(2) Pruning (line 7-23)
We present an illustration of Layer Collapse, as depicted in Figure 2. We begin by initializing the model $\mathcal{M}^{*}$ with the model $\mathcal{M}$ and set a layer pointer $l$ to start from $\mathcal{H}-\mathcal{C}$. Then, the iterative process of pruning begins:
RDSC Layer Merge (line 10-11) During each iteration, our approach involves merging the $\mathcal{K}$ layers following layer $l$ into layer $l$ itself and then discarding the redundant $\mathcal{K}$ layers, where $\mathcal{K}$ is the minimum of $\mathcal{C}-1$ and the total layer count of *$\mathcal{M}^{*}-l$, implying merging either the subsequent $\mathcal{C}-1$ layers or all layers following $l$, thus to prune the model $\mathcal{M}^{*}$, resulting in the interim model $\mathcal{M}_{tmp}$.
Calculate Similarity (line 12) We process each sentence in $\mathcal{D}$ using forward computations with $\mathcal{M}_{tmp}$ and $\mathcal{M}$ to derive their representations. For every sentence, we then calculate the cosine similarity between these representations from both models, averaging these values to obtain the overall similarity score $s$.
Merge Evaluation and Adjustment (line 13-21) Then, we evaluate $s$ against the threshold $\mathcal{T}$. Should $s$ exceed $\mathcal{T}$, the current merge is considered successful. Then, $\mathcal{M}_{tmp}$ is updated to $\mathcal{M}^{*}$ for the next iteration, and the pointer $l$ is adjusted downwards by $\mathcal{I}$ layers. Conversely, $l$ is simply reduced by a single layer. It is important to highlight that the instances may occur where $l$ falls below the total layer count of $\mathcal{M}^{*}$ after a series of successive merges. Consequently, it is required to reset $l$ to the layer count in *$\mathcal{M}^{*}-\mathcal{C}$, as illustrated in line 15.
@@ 2.3 Complexity Analysis
The complexity of LaCo primarily depends on model inference. In the worst-case scenario, where $\mathcal{L}$ is set to 0 and $\mathcal{H}$ is set to the total number of model layers. If in each iteration, the similarity $s$ is less than $\mathcal{T}$, then all layers will be traversed. Thus, the worst-case time complexity is $O(\mathcal{H}\times||\mathcal{D}||)$. For instance, considering Llama2-13B with 40 layers and $||\mathcal{D}||$ consisting of 10 sentences, the maximum number of inference steps would be only 400 sentences, which can be completed within minutes on a single GPU. Therefore, our approach demonstrates excellent time performance.
@@ 3.2 Benchmarks
To comprehensively evaluate the pruned model’s capabilities, we utilized the OpenCompass evaluation framework Contributors (2023). Specifically, following OpenCompass categorization, we conduct evaluations in five aspects: Reasoning, Language, Knowledge, Examination and Understanding. We select several benchmarks from each category. Reasoning: CMNLI Xu et al. (2020), HellaSwag (HeSw) Zellers et al. (2019), PIQA Bisk et al. (2019). Language: CHID Zheng et al. (2019), WSC Levesque et al. (2012). Knowledge: CommonSenseQA (CoQA) Talmor et al. (2018), BoolQ Clark et al. (2019). Examination: MMLU Hendrycks et al. (2021), CMMLU Li et al. (2023). Understanding: Race-High/Middle (H/M) Lai et al. (2017), XSum Narayan et al. (2018), C3 Sun et al. (2020).
We conduct evaluations using official scripts from OpenCompass, all zero-shot or few-shot, without additional training. Two evaluation modes are utilized: perplexity (PPL) and generation (GEN) 111opencompass.readthedocs.io/en/latest/get_started/faq.html. For CHID and XSum, we use the GEN mode. For the WSC dataset, we use both ${}_{\text{P}}$) and ${}_{\text{G}}$) modes. The remaining benchmarks are evaluated using the PPL mode. The evaluation results on each benchmark are converted to a score by OpenCompass, where a higher score indicates better performance. OpenCompass provides official evaluation results for the Baichuan2 and Llama2 series. However, to avoid discrepancies resulting from hardware and software environments, as well as potential errors in official results, we reproduce all results to ensure fairness.
@@ 3.3 Baselines
Since our LaCo involves structured pruning, which directly removes components from the LLM, we chose two current state-of-the-art (SOTA) structured pruning methods, LLM-Pruner (LLMPru.) Ma et al. (2023) and SliceGPT Ashkboos et al. (2024) which has surpassed the previous SOTA sparsity method, SparseGPT Frantar and Alistarh (2023) as our baselines. For experiments, we set the pruning ratio of these methods to be equivalent to or slightly smaller than our method to ensure fairness.
@@ 3.4 Settings
Table 1: The main results of our experiments. “Dense" is the official LLM results in OpenCompass and “$*{}^{*}$" is our reproduction. "LLMPru." and "SliceGPT" are two baseline comparisons.
@@ 3.5 Main Results
In Table 1, we present the results of four LLMs under different pruning methods across various benchmarks. “Dense” represents the official results of the unpruned LLMs in OpenCompass leaderboards, while “$*{}^{*}$” represents our reproduction of the “Dense” results. "LLMPru." and "SliceGPT" correspond to the two baselines, respectively. “Ratio" refers to the overall pruning ratio, namely the proportion of the total number of pruned parameters to that of the unpruned model. “Lay.” denotes the total number of layers in the model.
Comparing Dense and $*{}^{*}$, the results show not much difference, with most discrepancies within 5%. This indicates our experimental setup is error-free. To ensure fairness, we compare the results against $*{}^{*}$ in the subsequent analyses.
@@
To provide a more intuitive presentation of the results in Table 1, we compute the average scores of each pruner across all benchmarks (Avg.), the average scores per category (Reas., Lan., Know., Exam., Unde.), and the average performance percentages relative to $*{}^{*}$ across all benchmarks (Per.) in Table 2. Overall, our average scores are significantly higher than the baselines. It can also be inferred that Laco exhibits significantly superior performance overall in four out of five categories: Language, Knowledge, Examination, and Understanding. Even though there is a slight dip in performance in Reasoning, it remains comparable to the baselines. In addition, LaCo holds a significant advantage: relative to $*{}^{*}$, our average performance percentage across all datasets is far superior to the baselines. The average percentage surpasses 80% in three out of four models, with the lowest being on Baichuan2-7B, yet still exceeding 73%. However, none of the baselines have exceeded 70% on any model.
@@
Table 2: The average scores and the percentages comparison with the $*{}^{*}$.
To demonstrate the stability of models pruned by our LaCo method, we compute the percentage of result on each benchmark relative to $*{}^{*}$, as shown in Appendix C Table 7. We observe that models pruned by LaCo exhibit stability relative to the baselines, maintaining performance above 70% on most benchmarks. Furthermore, there were no instances of models crashing outright, with performance dropping below 30%.
[…]
In summary, our LaCo stands out as a superior pruning method, effectively preserving model performance and demonstrating exceptional stability across various benchmarks. Additionally, our approach is straightforward, relying solely on parameter differences and additions without necessitating modifications to the model’s internal structure, such as intermediate layer dimensions. This results in a more concise and efficient pruning solution.
4 Analysis

### v2

@@ Abstract
Large language models (LLMs) based on transformer are witnessing a notable trend of size expansion, which brings considerable costs to both model training and inference. However, existing methods such as model quantization, knowledge distillation, and model pruning are constrained by various issues, including hardware support limitations, the need for extensive training, and alterations to the model internal structure. In this paper, we propose a concise layer-wise structured pruner called Layer Collapse (LaCo), in which rear model layers collapse into a prior layer, enabling a rapid reduction in model size while preserving the model structure. Comprehensive experiments show that our method maintains an average task performance of over 80% at pruning ratios of 25-30%, significantly outperforming existing state-of-the-art structured pruning methods. We also conduct post-training experiments to confirm that the LaCo effectively inherits the parameters of the original model. Additionally, we perform ablation studies on various settings of LaCo. Finally, we discuss our motivation from the perspective of layer-wise similarity and evaluate the performance of the pruned LLMs across various pruning ratios [^1] https://github.com/yangyifei729/LaCo.
@@ 1 Introduction
Recently, large language models (LLMs) based on Transformer (Vaswani et al., 2017) have showcased impressive capabilities across diverse tasks. However, the prevailing trend in model development leans towards larger scales, placing substantial demands on computational resources.
To mitigate the above challenge, various approaches have been explored to reduce the inference and training costs of models or to derive a smaller model from an LLM, including model quantization Dettmers et al. (2022); Yao et al. (2022); Xiao et al. (2023), knowledge distillation Liu et al. (2022); Hsieh et al. (2023); Shridhar et al. (2023), and model pruning Zhang et al. (2022); Frantar and Alistarh (2023); Ma et al. (2023). However, existing solutions exhibit certain notable drawbacks. Model quantization typically necessitates specific hardware support and often impacts model performance. Knowledge distillation often requires retraining a smaller model, which is costly and task-specific. Model pruning, whether non-structured or structured, has its issues. Non-structured pruning often involves model sparsity, which generally leads to certain performance loss and also relies on hardware support. Structured pruning entails removing specific modules, often altering the model structure and diminishing the model portability.
@@
Specifically, we observe that merging the parameter differentials of certain layers with their subsequent layers often does not significantly impact model performance, as illustrated in Figure 1.
We term it the Reserving-Differences-while-Seeking-Common (RDSC) Layer Merge, as it incorporates parameter differencing and merging. Building upon this insight, we introduce a streamlined yet potent layer-wise pruner dubbed Layer Collapse (LaCo), in which rear layers collapse into a prior layer, with the objective of preserving the model’s output representation. In this paper:
@@
$\bullet$ The Layer Collapse preserves the internal structure of LLMs, such as maintaining intermediate dimensions. So, the pruned models can be seamlessly integrated into existing applications without any changes to the system’s implementation.
$\bullet$ We conduct post-training to confirm that Layer Collapse can efficiently inherit parameters and requires only minimal training to restore the pruned model to the original model’s loss convergence level. Additionally, we discuss our motivation and evaluate the performance of pruned models using LaCo across different pruning ratios. We also perform ablation studies on various settings of LaCo.
@@ 2.1 Reserving-Differences-while-Seeking-Common Layer Merge
For the $l$-th layer of an LLM, we denote all its parameters, including those in self-attention (SAN) and MLP as $\bm{\theta}_{l}$. For the $m$ consecutive layers following it, we merge the parameters of $\bm{\theta_{l+1}},\bm{\theta_{l+2}},\cdots,\bm{\theta_{l+m}}$ into $\bm{\theta_{l}}$ to form $\bm{\theta^{*}_{l}}$:
$\bm{\theta}^{*}_{l}$ $=\bm{\theta}_{l}+(\bm{\theta}_{l+1}-\bm{\theta}_{l})+\cdots+(\bm{\theta}_{l+m}-\bm{\theta}_{l})$ (1)
$=\bm{\theta}_{l}+\sum_{k=1}^{m}(\bm{\theta}_{l+k}-\bm{\theta}_{l})$
where $(\bm{\theta_{l+k}}-\bm{\theta_{l})}$ is the layer-wise parameter difference. Given identical layer structures, we independently apply these processes to both SAN and MLP. Then, these $m$ consecutive layers will be discarded. Subsequent model pruning will continuously involve RDSC Layer Merge which can be regarded as the continual collapse of layers onto specific layers, hence the name Layer Collapse.
@@
Algorithm 1 Workflow of Layer Collapse
1:LLM $\mathcal{M}$; Number of layers combined in each merge $\mathcal{C}$; Layer range $[\mathcal{L},\mathcal{H}]$; Minimum interval between two adjacent merged layers $\mathcal{I}$; Few-shot Calibration Samples $\mathcal{D}$; Threshold for representation similarity $\mathcal{T}$
2:Pruned LLM $\mathcal{M}^{*}$
@@
4:$l\leftarrow\mathcal{H}-\mathcal{C}$
5:while $l>=\mathcal{L}$ do
6:     $\mathcal{K}\leftarrow$ Min($\mathcal{C}-1$, $\mathcal{M}^{*}$)$-l$)
7:     $\mathcal{M}_{tmp}\leftarrow$ $\mathcal{M}^{*}$, $l$, $\mathcal{K}$)
8:     $s\leftarrow$ Avg_Cos_Sim($\mathcal{M}_{tmp}$, $\mathcal{M}$, $\mathcal{D}$)
9:     if $s>\mathcal{T}$ then
10:         $\mathcal{M}^{*}\leftarrow\mathcal{M}_{tmp}$
11:         $l\leftarrow l-\mathcal{I}$
12:         if $l>$ $\mathcal{M}^{*}$then
13:              $l\leftarrow$$\mathcal{M}^{*}$)$-\mathcal{C}$
14:         end if
15:     else
16:         $l\leftarrow l-1$
17:     end if
18:end while
19:return $\mathcal{M}^{*}$
@@ 2.2 Layer Collapse
We dynamically merge adjacent layers from the topmost layer down, ensuring the pruned model’s output representation on few-shot calibration samples remains as similar as possible to the original model to minimize performance loss. Algorithm 1 summarizes the workflow of Layer Collapse:
(1) Preparation
@@
(2) Pruning (line 1-17)
We present an illustration of Layer Collapse in Figure 2. We begin by initializing the model $\mathcal{M}^{*}$ with the model $\mathcal{M}$ and set a layer pointer $l$ to start from $\mathcal{H}-\mathcal{C}$. Then, the iterative process begins:
RDSC Layer Merge (line 4-5) During each iteration, our approach involves merging the $\mathcal{K}$ layers following layer $l$ into layer $l$ itself and then discarding the redundant $\mathcal{K}$ layers, where $\mathcal{K}$ is the minimum of $\mathcal{C}-1$ and the total layer count of $\mathcal{M}^{*}-l$, implying merging either the subsequent $\mathcal{C}-1$ layers or all layers following $l$, thus to prune the model $\mathcal{M}^{*}$, resulting in the interim model $\mathcal{M}_{tmp}$.
Calculate Similarity (line 6) We process each sentence in $\mathcal{D}$ using forward computations with $\mathcal{M}_{tmp}$ and $\mathcal{M}$ to derive their representations which are the output hidden-states of the last layer of the model. For every sentence, we then calculate the cosine similarity between these representations from both models, averaging these values to obtain the overall similarity score $s$.
Merge Evaluation and Adjustment (line 7-15) Then, we evaluate $s$ against the threshold $\mathcal{T}$. Should $s$ exceed $\mathcal{T}$, the current merge is considered successful. Then, $\mathcal{M}_{tmp}$ is updated to $\mathcal{M}^{*}$ for the next iteration, and the pointer $l$ is adjusted downwards by $\mathcal{I}$ layers. Conversely, $l$ is simply reduced by a single layer. It is important to highlight that the instances may occur where $l$ falls below the total layer count of $\mathcal{M}^{*}$ after a series of successive merges. Consequently, it is required to reset $l$ to the layer count in $\mathcal{M}^{*}-\mathcal{C}$, as illustrated in line 11.
@@ 2.3 Complexity Analysis
The complexity of LaCo primarily depends on model inference. In the worst-case scenario, with $\mathcal{L}$ set to 0 and $\mathcal{H}$ to the total number of layers, if in each iteration the similarity $s$ is less than $\mathcal{T}$, all layers will be traversed. Thus, the worst-case time complexity is $O(\mathcal{H}\times||\mathcal{D}||)$. For example, for Llama2-13B with 40 layers and $||\mathcal{D}||$ consisting of 10 sentences, the maximum number of inference steps would be only 400 sentences, which can be completed within minutes on a single GPU.
@@ 3.2 Benchmarks
To comprehensively evaluate the pruned model’s capabilities, we utilized the OpenCompass evaluation framework Contributors (2023). Specifically, following OpenCompass categorization, we conduct evaluations in five aspects: Reasoning, Language, Knowledge, Examination and Understanding. We select several benchmarks from each category. Reasoning: CMNLI Xu et al. (2020), HellaSwag (HeSw) Zellers et al. (2019), PIQA Bisk et al. (2019). Language: CHID Zheng et al. (2019), WSC Levesque et al. (2012). Knowledge: CommonSenseQA (CSQA) Talmor et al. (2018), BoolQ Clark et al. (2019). Examination: MMLU Hendrycks et al. (2021), CMMLU Li et al. (2023). Understanding: Race-High/Middle (H/M) Lai et al. (2017), XSum Narayan et al. (2018), C3 Sun et al. (2020).
We conduct evaluations using official scripts from OpenCompass, all zero-shot or few-shot, without additional training. Two evaluation modes are utilized: perplexity (PPL) and generation (GEN) 222opencompass.readthedocs.io/en/latest/get_started/faq.html. For CHID and XSum, we use the GEN mode. For the WSC dataset, we use both ${}_{\text{P}}$) and ${}_{\text{G}}$) modes. The remaining benchmarks are evaluated using the PPL mode. The evaluation results on each benchmark are converted to a score by OpenCompass, where a higher score indicates better performance. OpenCompass provides official evaluation results for the Baichuan2 and Llama2 series. However, to avoid discrepancies resulting from hardware and software environments, as well as potential errors in official results, we reproduce all results to ensure fairness.
@@ 3.3 Baselines
Since LaCo involves structured pruning, which directly removes components from LLMs, we select two state-of-the-art (SOTA) structured pruning methods, LLM-Pruner (LLMPru.) Ma et al. (2023) and SliceGPT Ashkboos et al. (2024), as our baselines. These methods have surpassed the previous SOTA sparsity method, SparseGPT Frantar and Alistarh (2023). In our experiments, we set the pruning ratios of baselines to be equivalent to or slightly smaller than LaCo to ensure fairness.
@@ 3.4 Settings
Table 1: The main results of our experiments. Lay. is the number of model layers. Dense is the official LLM results in OpenCompass and Dense is our reproduction. LLMPru. and SliceGPT are two baseline comparisons.
@@ 3.5 Main Results
In Table 1, we present the results of four LLMs under different pruning methods across various benchmarks. “Dense” represents the official results of the unpruned LLMs in OpenCompass leaderboards, while “Dense” represents our reproduction of the “Dense” results. "LLMPru." and "SliceGPT" correspond to the two baselines, respectively. “Ratio" refers to the overall pruning ratio, namely the proportion of the total number of pruned parameters to that of the unpruned model. “Lay.” denotes the total number of layers in the model.
Comparing Dense and Dense, the results show not much difference, with most discrepancies within 5%. This indicates our experimental setup is error-free. To ensure fairness, we compare the results against Dense in the subsequent analyses.
@@
To provide a more intuitive presentation of the results in Table 1, we compute the average scores of each pruner across all benchmarks (Avg.), the average scores per category (Reas., Lan., Know., Exam., Unde.), and the average performance percentages relative to Dense across all benchmarks (Per.) in Table 2. Overall, our average scores are significantly higher than the baselines. LaCo shows superior performance in four out of five categories. Though there is a slight dip in Reasoning, it remains comparable. Additionally, LaCo’s average performance percentage across all datasets, relative to Dense, is far superior to the baselines. The average percentage surpasses 80% in three out of four models, with the lowest being 73% on Baichuan2-7B. In contrast, none of the baselines exceed 70%.
To demonstrate the stability of the pruned models by LaCo, we compute the performance percentage relative to Dense (Appendix D.4 Table 22). LaCo-pruned models maintain performance above 70% on most benchmarks and do not experience crashes, with no performance dropping below 30%.
Notably, on three benchmarks evaluated through GEN mode, CHID, XSUM, and WSCG, the LLMs pruned by LaCo maintain relatively stable performance, while models pruned by baselines exhibit poorly, with even multiple results becoming 0.00. GEN mode scores are based on the model’s generated sentences, and the models pruned by baselines are prone to producing meaningless repetitive outputs. In Appendix D.5 Table 23, we showcase an example from the Xsum benchmark, where Llama2-7B, pruned by baselines, produces nonsensical repeated outputs, whereas our LaCo yields outputs resembling normal sentences.
We also conduct experiments with Llama2-70B on several benchmarks. The results in Appendix D.1 Table 19 show that LaCo still outperforms the baseline on larger-scale model.
In summary, LaCo is a superior pruner that preserves model performance and demonstrates exceptional stability across various benchmarks. It relies solely on parameter differences and additions, without altering the model’s internal structure, resulting in a concise and efficient pruning solution.
@@
Table 2: The average scores and the percentages comparison with the Dense.
3.6 Comparison of Perplexity
Pruner Dense LaCo LLM-Pruner SliceGPT
PPL 4.46 13.93 17.30 14.51
Table 3: PPL for different pruners.
Since PPL is also a commonly used metric for evaluating model performance, we want to compare how LaCo differs from the baseline in maintaining the model’s PPL. We evaluate the PPL of the Llama2-7B model with 27% sparsity using different pruners. The evaluation is performed on 500 sentences selected from Wikipedia, each with a fixed length of 512 tokens.
The results in Table 3 show that the model pruned by LaCo also achieves a lower PPL compared to other baselines, further highlighting the advantage of our LaCo.
3.7 Pruning Time
Pruner LaCo LLM-Pruner SliceGPT
Pruning Time 14.7s 15.9s 313s
Table 4: Pruning time for different pruners.
To verify that LaCo has lower time complexity and faster pruning speed than the baselines, we compare LaCo with them for 27% sparsity pruning of Llama2-7B on a single A100 GPU. For fairness, we only measure the main pruning process, excluding the time for loading models, loading data, and storing models. The results in Table 4 show LaCo pruning is more efficient compared to the baselines.
3.8 Memory Usage and Inference Speed
We also aim to investigate whether the model pruned by LaCo offers advantages in memory usage and inference speed compared to the models pruned by the baselines. In Table 5, we present the average memory consumption and inference speed of the Llama2-13B pruned models from Table 1 on the English Wiki dataset (The results for all models are in Appendix D.2 Table 20). All models are loaded in Bf16 on a single A100 GPU.
Pruner LaCo Dense LLMPru. SliceGPT
Memory 19422 25902 19874 22506
Infer. 38.65 29.98 27.15 ($\downarrow$) 35.16
Table 5: Memory usage (MB) and inference speed (tokens/s) of the Llama2-13B pruned by different pruners. $\downarrow$ indicates performance worse than the Dense model.
The results indicate that the LaCo-pruned models consume less memory and achieve faster inference speeds. Moreover, while existing baselines may decrease inference speeds compared to the dense model, LaCo does not have this issue.
4 Further Analysis